a particle confined to motion along x axis moves with a constant accelaration fro x=2m to x=8m during the 2.5secs time interval. the velocity of particle at x=8m is 2.8m/s. what is the accelaration during this time interval?

Grade:12

1 Answers

Swapnil Saxena
12 years ago

Make use of Newtons 1, 3  equation

t=2.5, s=6, v=2.8

v= u+a(2.5) --(i)

v^2= u^2+2a(6)---(iii)

Using equation I and II we get (v+u )=24/5= 4.8

When v= 2.8 the the initial velocity of the particle is 2m/s

Usin i equation

Acceleartion is  0.8/2.5 = 8/25 = 0.32m/s^2